# RedDebate: Safer Responses through Multi-Agent Red Teaming Debates

### v1

@@ Abstract
We propose RedDebate, a novel multi-agent debate framework that leverages adversarial argumentation among Large Language Models (LLMs) to proactively identify and mitigate their own unsafe behaviours. Existing AI safety methods often depend heavily on costly human evaluations or isolated single-model assessment, both subject to scalability constraints and oversight risks. RedDebate instead embraces collaborative disagreement, enabling multiple LLMs to critically examine one another’s reasoning, and systematically uncovering unsafe blind spots through automated red-teaming, and iteratively improve their responses. We further integrate distinct types of long-term memory that retain learned safety insights from debate interactions. Evaluating on established safety benchmarks such as HarmBench, we demonstrate the proposed method’s effectiveness. Debate alone can reduce unsafe behaviours by 17.7%, and when combined with long-term memory modules, achieves reductions exceeding 23.5%. To our knowledge, RedDebate constitutes the first fully automated framework that combines multi-agent debates with red-teaming to progressively enhance AI safety without direct human intervention. [^1] https://github.com/aliasad059/RedDebate
RedDebate: Safer Responses through Multi-Agent Red Teaming Debates
Ali Asad, Stephen Obadinma, Radin Shayanfar, Xiaodan Zhu Department of Electrical and Computer Engineering & Ingenuity Labs Research Institute Queen’s University, Kingston, Canada {ali.asad, 16sco, radin.shayanfar, xiaodan.zhu}@queensu.ca
@@ 1 Introduction
Humans have long employed debate and dialogue as powerful methods to refine their understanding and correct misconceptions, a tradition deeply rooted in the Socratic method, or Elenchus Seeskin (1987); Paul and Elder (2008). Recently, advances in Artificial Intelligence (AI), specifically in Large Language Models (LLMs) have equipped machines with unprecedented capabilities for natural language communication. These advances include enhanced understanding, reasoning, and generating of language grounded in extensive accumulated knowledge (Minaee et al., 2024; OpenAI, 2024). These foundational skills provide an opportunity to explore debate as a means to leverage the inherent strengths of these models (Bench-Capon and Dunne, 2007). Specifically, debate can facilitate examining complex ideas, evaluating evidence, and critically analyzing competing viewpoints. As a result, there is growing interest in using debate to enhance and harness the strengths of LLMs (Irving et al., 2018; Parrish et al., 2022; Du et al., 2024).
LLM-based debate explicitly leverages the diversity of perspectives present within different LLMs, allowing these highly capable language models to contribute unique insights within structured dialogues (Liang et al., 2024). Therefore, this richer collaborative process has the potential to enhance the depth, factuality, and robustness of model-generated reasoning (Du et al., 2024; Khan et al., 2024). Indeed, debate has been shown to effectively surface flaws in reasoning, proving particularly valuable for enabling human judges or comparatively less capable AI systems to detect incorrect or potentially unsafe reasoning from superhuman-level AI (Irving et al., 2018; Khan et al., 2024). Thus, debating represents a promising approach to address key safety concerns within AI systems by effectively tapping into the advanced capabilities of modern LLMs and fostering collective reasoning.
@@
Figure 1: RedDebate framework overview. Agents debate a red-teaming prompt over multiple rounds, refining responses through peer interaction. An evaluator performs a further safety check, and key safety insights are then generated and stored in memory for future use, enabling continuous safety improvement.
@@
One prominent challenge in the AI safety domains is the detection and mitigation of unsafe AI responses—commonly addressed through red-teaming methods. These unsafe responses can be characterized as arising from the AI system’s misconceptions regarding alignment with human values. Traditional approaches to red-teaming and alignment heavily depend on human annotations and evaluations, facing challenges in scalability, due to critical oversights, and incur significant costs, as LLMs continue to grow in complexity and size (Bowman et al., 2022; Perez et al., 2022; Bai et al., 2022c). AI-driven supervision methods present an alternative by automating these processes through the enhanced reasoning capabilities of LLMs. Most of these approaches rely on single-agent supervision, such as constitutional AI using self-criticism under predefined constitutions (Bai et al., 2022c) and self-reflection (Shinn et al., 2023). The key issue with these methods is that an agent might not realize its own mistakes or knowledge gaps, making it difficult to reflect and correct itself effectively (Liang et al., 2024).
@@
To address these limitations, we propose RedDebate, a novel multi-agent debate framework designed to enhance AI safety by enabling collaborative, adversarial exploration of unsafe reasoning, fostering deeper insights unattainable through individual self-assessment alone. Our framework combines the collective reasoning strengths inherent in debate, including diversity of perspectives and mutual critical assessment, along with systematic vulnerability detection through automated red-teaming. With this objective in mind, we explore the central research question: Can a set of LLMs collaboratively identify and refine their unsafe behaviours through debate?
@@
  • We introduce variant types of long-term memory tested across diverse debate scenarios, showing that memory significantly enhances agent performance via dynamic feedback retention and retrieval, while debates expose distinct vulnerabilities.
  • We demonstrate that LLMs effectively learn from debate and memory to reduce unsafe responses, achieving up to 17.7% reduction with debate alone and over 23.5% when combined with long-term memory on HarmBench.
@@ 2 Related Work
Red Teaming
exposes unsafe model behaviour via adversarial prompts. While early work relied on human-written tests, recent methods automate attack generation and evaluation using LLMs (Perez et al., 2022), with extensions for broader coverage (Hong et al., 2024; Casper et al., 2023). Our work advances this by creating a fully automated red-teaming pipeline where LLM agents generate, critique, and evaluate responses.
@@
Feedback, either human or model-generated, helps steer LLM behaviour. RLHF finetunes models based on human preferences (OpenAI et al., 2024; Bai et al., 2022a), while Bai et al. (2022b) and Shinn et al. (2023) use self-generated feedback for response improvement via predefined rules or textual feedback derived in subsequent trials. We extend this with peer critique in debate to produce richer feedback, combined with long-term memory for enhanced retention.
@@
allows determining if and how some actions could be enforced to increase safety in a system (Dong et al., 2024b; Rebedea et al., 2023; Guardrails AI, ). NeMo Guardrails (Rebedea et al., 2023) uses Colang [^2] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html to define safe and highly flexible conversational flows. Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. We build on this by introducing guardrails as a form of long-term memory for safety learning. Please refer to Appendix A for a more detailed overview of related works for each area.
@@ 3.1 RedDebate Framework
As illustrated in Figure 2, the process begins by selecting a red-teaming prompt $\mathcal{X}\in\mathcal{P}$ from a set of adversarial prompts $\mathcal{P}$, which serves as the central topic for the debate. A set of $N$ debater agents $\mathcal{D}=\{\mathrm{D_{1}},\mathrm{D_{2}},\ldots,\mathrm{D_{N}}\}$ simultaneously generate responses at each round $t$, as defined in Equation 1,
@@
where $\mathcal{M}_{\text{S}}$ and $\mathcal{M}_{\text{L}}$ denote the shared short-term and long-term memory, respectively. At each round $t$, the set of agents responses $\mathcal{R}^{(t)}=\{r_{n}^{(t)}\}_{n=1}^{N}$ will be stored in $\mathcal{M}_{\text{S}}$. This enables each agent to access not only its own previous answers but also those of other agents. In subsequent rounds, agents use this memory to critique others’ responses, refine their previous statements, or offer feedback to peers. After a fixed number of rounds of debate $T$, a separate evaluator agent $\mathrm{E}$ assesses the safety of all responses =$\mathcal{R}=\{r_{n}^{(t)}\}_{n=1,t=1}^{N,T}$ generated throughout the debate and produces binary labels $\mathcal{Y}=\{y_{n}^{(t)}\}_{n=1,t=1}^{N,T}$, as described by Equation 2, where each label $y_{n}^{(t)}\in\{0,1\}$ indicates whether the response is safe.
@@
$\mathcal{M}_{\text{S}}\leftarrow\emptyset$
  // Initialize STM
$\mathcal{R}\leftarrow[]$
  // Initialize Debate History
@@ 8             $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{r_{n}^{(t)}\}$
       $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathcal{R}^{(t)}\}$
        // Append to History
12
@@
  // Return Debate History
Algorithm 1 Multi-Agent Debate with Memory
@@ 3.2 Debate Strategies
In PReD, often responses generated by debating agents can overlap or represent similar reasoning, reducing the diversity and critical scope of the discussion. To address this limitation, we explore the introduction of auxiliary agents that intentionally diversify the debate landscape. These agents enhance robustness by exposing the debating agent to explicitly contrasting perspectives, prompting a more critical reassessment of the original response.
To achieve our goal—and inspired by Liang et al. (2024) who promote perspective diversification, though in a different context—we integrate two auxiliary agents in the safety setting: the Devil agent ($\mathrm{D}^{\ominus}$) and the Angel agent ($\mathrm{D}^{\oplus}$). When a debating agent produces a round-$t$ response $r_{n}^{(t)}$, the Devil agent generates a rejection $\delta_{n}^{(t)}=\mathrm{D}^{\ominus}(r_{n}^{(t)})$, which critically challenges the response, acting as a skeptical adversary independent of the actual safety status. Conversely, the Angel agent provides supportive reinforcement $\alpha_{n}^{(t)}=\mathrm{D}^{\oplus}(r_{n}^{(t)})$, explicitly affirming and encouraging the original response. With these explicitly contrasting viewpoints—one inherently skeptical (possibly unsafe) and the other explicitly supportive (likely perceived as safe)—the agent is compelled to critically reassess its reasoning. This opposing feedback is anticipated to facilitate self-refinement and correction of potential unsafe responses. If the agent does not adequately reflect or becomes misled by persuasive auxiliary input, such inadequate responses are more reliably identified and flagged by the evaluator $\mathrm{E}$ and feedback generator $\mathrm{F}$ for further corrective intervention.
@@
We observe that agents often do not proactively provide feedback or refine their responses based on peer contributions unless explicitly triggered or challenged by others. Additionally, assessing the depth and stability of an agent’s conviction in its responses—whether these are firmly held or merely superficial positions subject to change—proved crucial for effective deliberation. To address these concerns, we introduce a novel scenario incorporating an auxiliary Socratic agent $\mathrm{D^{S}}$ alongside the basic debaters $\mathcal{D}$. Within this setting, the Socratic agent assumes a distinct questioning role, critically examining the responses $\mathcal{R}^{(t)}$ provided at debate round $t$. Its primary objective is to uncover implicit assumptions, highlight reasoning gaps, and promote deeper reflection by actively requesting evidence, illustrative examples, or clarification as needed. Inspired by traditional Socratic dialogues Paul and Elder (2008); Chang (2023), this approach systematically probes the coherence, depth, and soundness of agent arguments.
Pseudo-codes for the DAReD and SReD strategies are provided in Appendix G.
@@ 3.3 Memory
Memory is crucial for intelligent reasoning, enabling agents to learn from past experiences, refine decision-making, and avoid repeating errors. In the context of RedDebate, where agents engage iteratively in challenging debate, effective memory mechanisms allow continuous enhancement of agent behaviour. Inspired by cognitive structures in human decision-making Zhang et al. (2024), we propose to integrate two complementary memory modules into RedDebate: short-term memory (STM) and long-term memory (LTM). STM provides immediate context, maintaining coherence within an ongoing debate, and is reset upon each debate’s completion. LTM, on the other hand, acts as a persistent repository, storing accumulated safety insights and feedback. In this section, we explore four variations of LTM tailored to our setting.
@@
is widely adopted in prior work due to its interpretability, ease of implementation, and efficient read-write operations Zhang et al. (2024). In this memory type, generated natural language feedback is incorporated into the agent’s prompt to make the agent aware of previously learned lessons. However, since agents may accumulate a large number of feedback entries through repeated trial and error, including all of them in the prompt can be inefficient, given our setup with relatively lightweight LLMs and limited context windows. To address this, we encode all feedback $\Phi=\{\phi_{1},\ldots,\phi_{k}\}\subseteq\mathcal{M}_{L}$ into vector representations [^3] Vector representations are used only for retrieval; ultimately, agents receive top-matching textual feedback entries. and store them in a vector database. For future prompts $\mathcal{X}^{\prime}$, the system retrieves the most relevant feedback entries using a similarity function $\mathrm{sim}(\mathcal{X}^{\prime},\phi)$ and adds these textual feedbacks to the agent’s context. This ensures memory remains concise, relevant, and helpful without overwhelming the agent.
@@
aims to explicitly encode unsafe experiences into executable programmatic constraints or guardrails. A core limitation of prior types of memory is that they are inherently passive: agents must actively integrate lessons into generation. Consequently, the burden resides on the agent to correctly recall and interpret memory content at generation. Inspired by recent work on automatic programmatic guardrail generation (Sreedhar et al., 2024; Dong et al., 2024a), we adapt the idea to the safety setting by implementing GLTM to automatically encode agents’ past unsafe experiences as guardrails. Before an agent generates a response, the input prompt will be directly rejected if it matches a previously known unsafe flow.
We use one-shot prompting to generate guardrails given a generated feedback $\phi$ and the corresponding harmful question $\mathcal{X}$. We instruct the LLM to output an expression and a list of examples for each $(\mathcal{X},\phi)$ pair. The expression is an abstraction of the harmful user behaviour, similar to abstractions generated in Wang et al. (2024b). It also serves as the flow names–Colang equivalent of functions–in the final resulting Colang application. Additionally, we instruct the LLM to provide examples of user utterances of such harmful behaviours. We merge the examples of the expressions if the expressions are the same. Finally, we use a rule-based method to convert the expressions into Colang flows, allowing use of Colang’s built-in intent detection feature to match the defined harmful behaviours and reject the user request. Figure 19 in Appendix shows an example of the code generation prompt and the resulting guardrail flow.

### v2

@@ Abstract
We introduce RedDebate, a novel multi-agent debate framework that provides the foundation for Large Language Models (LLMs) to identify and mitigate their unsafe behaviours. Existing AI safety approaches often rely on costly human evaluation or isolated single-model assessment, both constrained by scalability and prone to oversight failures. RedDebate employs collaborative argumentation among multiple LLMs across diverse debate scenarios, enabling them to critically evaluate one another’s reasoning and systematically uncover unsafe failure modes through fully automated red-teaming. We further integrate distinct long-term memory modules that preserve safety-relevant insights from debate interactions and leverage them during subsequent inference, facilitating continuous refinement of model behaviour. Empirical evaluation on safety benchmarks across a diverse set of models demonstrates that RedDebate substantially reduces unsafe outputs. While debate alone allows LLMs to refine their behaviour, the addition of memory yields further significant reductions. To the best of our knowledge, RedDebate is the first fully automated framework to unify multi-agent debate and red-teaming to progressively enhance LLM safety without human intervention.  [^1] https://github.com/aliasad059/RedDebate
@@ 1 Introduction
Humans have long employed debate and dialogue as powerful methods to refine their understanding and correct misconceptions, a tradition deeply rooted in the Socratic method, or Elenchus Seeskin (1987); Paul & Elder (2008). Recent advances in Artificial Intelligence (AI), specifically in Large Language Models (LLMs), have demonstrated unprecedented capabilities to comprehend, reason, and generate language grounded in vast knowledge corpora (Minaee et al., 2024; OpenAI, 2024). These foundational skills make debate a natural mechanism for leveraging LLMs; rather than treating a model’s output as a final answer, debate frames it as a claim to be tested, challenged, and iteratively improved. This perspective has driven growing interest in using debate to amplify the strengths of LLMs and mitigate their weaknesses (Irving et al., 2018; Parrish et al., 2022; Du et al., 2024). By enabling models to present, defend, and revise their reasoning, debate can reveal hidden assumptions, surface counterexamples, and drive convergence toward more robust conclusions (Bench-Capon & Dunne, 2007). Moreover, Multi-Agent Debate (MAD) explicitly leverages the diversity of perspectives present within different LLM agents, allowing these highly capable language models to contribute unique insights within structured dialogues (Liang et al., 2024). Therefore, this richer collaborative process has the potential to enhance the depth, factuality, and robustness of model-generated reasoning (Du et al., 2024; Khan et al., 2024). Indeed, debate has been shown to effectively surface flaws in reasoning, proving particularly valuable for enabling human judges or comparatively less capable AI systems to detect incorrect or potentially unsafe reasoning from superhuman-level AI (Irving et al., 2018; Khan et al., 2024).
@@
Figure 1: RedDebate framework overview. Multiple agents debate a red-teaming prompt across several rounds, refining their responses through peer interaction. An evaluator analyzes the outputs, flags unsafe patterns, and provides feedback. Distilled safety insights are stored in memory to prevent similar mistakes, enabling continual automated improvement.
@@
One prominent challenge in AI safety is the detection and mitigation of unsafe AI responses, commonly addressed through red-teaming methods, which often stem from the model’s misconceptions about human values or safety requirements. These misconceptions can manifest as flawed reasoning or unsafe behaviour in response to red-teaming prompts. Traditional approaches to uncovering such failures rely heavily on human-driven red-teaming and manual evaluation, processes that struggle to scale, risk missing subtle errors, and incur significant costs as models grow in complexity and size (Bowman et al., 2022; Perez et al., 2022; Bai et al., 2022b). This motivates the development of automated methods that not only detect unsafe behaviour but also help models recognize and correct their misconceptions. In the long term, such approaches could provide a practical mechanism for aligning and controlling increasingly capable AI systems, even at superhuman levels, by forcing them to critically examine and justify their reasoning before acting (Dalrymple et al., 2024; Tegmark & Omohundro, 2023). AI-driven supervision methods present an alternative by automating these processes through the enhanced reasoning capabilities of LLMs. Most of these approaches rely on single-agent supervision, such as constitutional AI using self-criticism under predefined constitutions (Bai et al., 2022b) and self-reflection (Shinn et al., 2023). The key limitation with these methods is that an agent might not realize its own mistakes or knowledge gaps, making it difficult to reflect and correct itself effectively (Liang et al., 2024).
@@
To address these limitations, we propose RedDebate, a novel multi-agent debate framework designed to enhance AI safety by enabling collaborative, adversarial exploration of unsafe reasoning, fostering deeper insights unattainable through individual self-assessment alone. Our framework combines the collective reasoning strengths inherent in debate, including diversity of perspectives and mutual critical assessment, along with systematic vulnerability detection through automated red-teaming. With this objective in mind, we explore the central research question: “Can a set of LLMs collaboratively identify and refine their unsafe behaviours through debate?”
@@
  • We introduce variant types of long-term memory tested across diverse debate scenarios and models, showing that memory significantly enhances agent safety performance via dynamic feedback retention and retrieval, while debates expose distinct vulnerabilities.
  • We demonstrate that LLMs effectively learn from debate to reduce unsafe responses, achieving up to 17.7% reduction with debate alone and over 23.5% when combined with long-term memory on HarmBench.
@@ 2 Related Work
Red-Teaming
exposes unsafe model behaviour via adversarial prompts. While early work relied on human-written tests, recent methods automate attack generation and evaluation using LLMs (Perez et al., 2022; Ge et al., 2023; Samvelyan et al., 2025), with extensions for broader coverage (Hong et al., 2024; Casper et al., 2023). Our work advances this by creating a fully automated red-teaming pipeline where a group of LLM agents generate, critique, and evaluate responses.
@@
Feedback, either human or model-generated, helps steer LLM behaviour. RLHF finetunes models based on human preferences (OpenAI et al., 2024; Bai et al., 2022a), while Bai et al. (2022b) and Shinn et al. (2023) use self-generated feedback for response improvement via predefined rules or textual feedback derived in subsequent trials. We extend this with peer critique in debate to produce richer feedback, combined with various long-term memory modules for enhanced retention.
@@
allows determining if and how some actions could be enforced to increase safety in a system (Dong et al., 2024b; Rebedea et al., 2023; Guardrails AI, 2025). NeMo Guardrails (Rebedea et al., 2023) uses Colang to define safe and highly flexible conversational flows. Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. We build on this by introducing guardrails as a form of long-term memory for safety learning. Please refer to Appendix A for a more detailed overview of related works.
@@ 3.1 RedDebate Framework
As illustrated in Figure 2, the process begins by selecting a red-teaming prompt $\mathcal{X}\in\mathcal{P}$ from a set of adversarial prompts $\mathcal{P}$, which serves as the central topic of the debate. A set of $N$ debater agents, $\mathcal{D}=\{\mathrm{D_{1}},\mathrm{D_{2}},\ldots,\mathrm{D_{N}}\}$ simultaneously generate responses $r_{n}^{(t)}$ in each round $t$, as defined in Equation 1,
@@
where $\mathcal{M}_{\text{S}}$ and $\mathcal{M}_{\text{L}}$ denote the shared short-term and long-term memory, respectively. At each round $t$, the set of agents responses $\mathcal{R}^{(t)}=\{r_{n}^{(t)}\}_{n=1}^{N}$ will be stored in $\mathcal{M}_{\text{S}}$. This enables each agent to access not only its own previous answers but also those of other agents. In subsequent rounds, agents use this memory to critique others’ responses, refine their previous statements, or offer feedback to peers. After a fixed number of rounds of debate $T$, a separate evaluator agent $\mathrm{E}$ assesses the safety of all responses $\mathcal{R}=\{r_{n}^{(t)}\}_{n=1,t=1}^{N,T}$ generated throughout the debate and produces binary labels $\mathcal{Y}=\{y_{n}^{(t)}\}_{n=1,t=1}^{N,T}$, as described by Equation 2, where each label $y_{n}^{(t)}\in\{0,1\}$ indicates whether the response is safe.
@@
$\mathcal{M}_{\text{S}}\leftarrow\emptyset,\mathcal{R}\leftarrow[]$
// Initiatation
@@ 8    $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{r_{n}^{(t)}\}$
12 $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathcal{R}^{(t)}\}$
@@
// Return History
20
Algorithm 1 Multi-Agent Debate
@@ 3.2 Debate Strategies
In PReD, often responses generated by debating agents can overlap or represent similar reasoning, reducing the diversity and critical scope of the discussion. To address this limitation, we explore the introduction of auxiliary agents that intentionally diversify the debate landscape. These agents enhance robustness by exposing the debating agent to explicitly contrasting perspectives, prompting a more critical reassessment of the original response. Inspired by Liang et al. (2024) who promote perspective diversification, though in a different context, we integrate two auxiliary agents in the safety setting: the Devil agent ($\mathrm{D}^{\ominus}$) and the Angel agent ($\mathrm{D}^{\oplus}$). When a debating agent produces a round-$t$ response $r_{n}^{(t)}$, the Devil agent generates a rejection $\delta_{n}^{(t)}=\mathrm{D}^{\ominus}(r_{n}^{(t)})$, which critically challenges the response, acting as a skeptical adversary independent of the actual safety status. Conversely, the Angel agent provides supportive reinforcement $\alpha_{n}^{(t)}=\mathrm{D}^{\oplus}(r_{n}^{(t)})$, explicitly affirming and encouraging the original response. With these explicitly contrasting viewpoints, one inherently skeptical and the other explicitly supportive (likely perceived as safe), the agent is compelled to critically reassess its reasoning.
@@
We observe that agents often do not proactively provide feedback or refine their responses based on peer contributions unless explicitly triggered or challenged by others. Additionally, assessing the depth and stability of an agent’s conviction in its responses, whether these are firmly held or merely superficial positions subject to change, proved crucial for effective deliberation. To address these concerns, we introduce a novel scenario incorporating an auxiliary Socratic agent $\mathrm{D^{S}}$ alongside the basic debaters $\mathcal{D}$. Within this setting, the Socratic agent assumes a distinct questioning role, critically examining the responses $\mathcal{R}^{(t)}$ provided at debate round $t$. Its primary objective is to uncover implicit assumptions, highlight reasoning gaps, and promote deeper reflection by actively requesting evidence, illustrative examples, or clarification as needed. Inspired by traditional Socratic dialogues Paul & Elder (2008); Chang (2023), this approach systematically probes the coherence, depth, and soundness of agent arguments. Pseudo-codes for the DAReD and SReD strategies are provided in Appendix D.
@@ 3.3 Memory
Memory is crucial for intelligent reasoning, enabling agents to learn from past experiences, refine decision-making, and avoid repeating errors. In the context of RedDebate, where agents engage iteratively in challenging debate, effective memory mechanisms allow continuous enhancement of agent behaviour. Inspired by cognitive structures in human decision-making Zhang et al. (2024), we propose to integrate two complementary memory modules into RedDebate: Short-Term Memory (STM, $\mathcal{M}_{\text{S}}$) and Long-Term Memory (LTM, $\mathcal{M}_{\text{L}}$). STM provides immediate context, maintaining coherence within an ongoing debate, and is reset upon each debate’s completion. LTM, on the other hand, acts as a persistent repository, storing accumulated safety insights and feedback. We explore four variations of LTM tailored to our setting in this section.
@@
is widely adopted in prior work due to its interpretability, ease of implementation, and efficient read-write operations Zhang et al. (2024). In this memory type, generated natural language feedback is incorporated into the agent’s prompt to make the agent aware of previously learned lessons. However, since agents may accumulate a large number of feedback entries through repeated trial and error, including all of them in the prompt can be inefficient, given our setup with relatively lightweight LLMs and limited context windows. To address this, we encode all feedback $\Phi=\{\phi_{1},\ldots,\phi_{k}\}\subseteq\mathcal{M}_{L}$ into vector representations and store them in a vector database. For future prompts $\mathcal{X}^{\prime}$, the system retrieves the most relevant feedback vectors using a similarity function $\mathrm{sim}(\mathcal{X}^{\prime},\phi)$ and adds their corresponding textual feedback to the agent’s context. This ensures memory remains concise, relevant, and helpful without overwhelming the agent.
@@
aims to explicitly encode unsafe experiences into executable programmatic constraints or guardrails. A core limitation of prior types of memory is that they are inherently passive: agents must actively integrate lessons into generation. Consequently, the burden resides on the agent to correctly recall and interpret memory content at generation. Inspired by recent work on automatic programmatic guardrail generation (Sreedhar et al., 2024; Dong et al., 2024a), we adapt the idea to the safety setting by implementing GLTM to automatically encode agents’ past unsafe experiences as guardrails. Before an agent generates a response, the input prompt will be directly rejected if it matches a previously known unsafe flow. We use one-shot prompting to generate guardrails given a generated feedback $\phi$ and the corresponding red-teaming prompt $\mathcal{X}$. The LLM is instructed to output both an expression and a list of examples for each $(\mathcal{X},\phi)$ pair. The expression abstracts the harmful user behavior, similar to those generated in Wang et al. (2024b), and also serves as the flow name—Colang’s equivalent of a function—in the resulting Colang application. Additionally, we instruct the LLM to provide examples of user utterances exhibiting the harmful behaviour. If multiple expressions are identical, we merge their examples. Finally, we apply a rule-based method to convert the expressions into Colang flows, leveraging Colang’s built-in intent detection feature to match the defined harmful behaviours and reject the user request. Figure 22 in the Appendix illustrates the code generation prompt and the resulting guardrail flow.